# Pinched geometries in $\mathbf{2}$D Lorentzian quantum Regge calculus

### v1

@@ 2.1 $\mathbf{2}$D Lorentzian QRC
where $x=\tau_{1}^{2}$, $y=\tau_{2}^{2}$ and $z=\sigma^{2}$. The Euclidean case requires the triangle inequalities to form a regular triangle. However, such extra constraints are not necessary in the 22D Lorentzian case because $A(x,y,z)$ is positive for given three edges satisfying $x,y\geq 0$ and $z>0$.
Let $N$ be the number of triangles and $N_{e}$ be the number of edges in the given triangulation. Note that $N$ is an even number, $N_{e}$ is given by $N_{e}=3N/2$, the number of space-like edges is $N/2$ and that of time-like edges is $N$. We then label each triangle and let $A_{s}$ be the area of the $s$-th triangle where $s=1,2.\cdots,N$. We also label the edges whose lengths are given by $\ell_{n}\,(n=1,2,\cdots,N_{e})$.
@@
As discussed in Ref. [52], we can perform an analytic continuation to change $iS$ of Eq. (2.3) to S-S while keeping the definitions of the triangle area (2.1) and the integral measures (2.4), in this model. We may consider the following partition function instead of Eq. (2.3):
@@ 2.2 Spike and pinched geometry
In our previous paper [52], we numerically conformed the suppression of pinched geometries through the finiteness of $\left<\tau^{2}\right>_{A}$ for a regular triangulation shown in Fig. 1. For different triangulations and integral measures, we verify the suppression by examining the higher moments $\left<\tau^{n}\right>_{A}$ up to $n=10$ in Sec. 3.
@@ 3.1 Discretization of tensor indices
Let $Z^{\prime}$ with insertion of an operator $\cal O$} be $I_{{\cal O}}$. The expectation value of $\cal O$} is then given by a ratio $I_{{\cal O}}/Z^{\prime}$. If $\cal O$} is local, then $I_{{\cal O}}$ can also be expressed as a tensor network with an impurity tensor corresponding to the operator insertion. The expectation value is thus evaluated as a ratio of two tensor networks.

### v2

@@ 2.1 $\mathbf{2}$D Lorentzian QRC
where $x=\tau_{1}^{2}$, $y=\tau_{2}^{2}$ and $z=\sigma^{2}$, and the inner products between two vectors are defined with respect to the Minkowski metric, $\eta=\text{diag}(-1,1)$. The Euclidean case requires the triangle inequalities to form a regular triangle. However, such extra constraints are not necessary in the 22D Lorentzian case because $A(x,y,z)$ is positive for given three edges satisfying $x,y\geq 0$ and $z>0$.
Let $N$ be the number of triangles and $N_{e}$ be the number of edges in the given triangulation. Note that $N$ is an even number, $N_{e}$ is given by $N_{e}=3N/2$, the number of space-like edges is $N/2$ and that of time-like edges is $N$. We then label each triangle and let $A_{s}$ be the area of the $s$-th triangle where $s=1,2.\cdots,N$. We also label the edges whose lengths are given by $\ell_{e}\,(e=1,2,\cdots,N_{e})$.
@@
As discussed in Ref. [29], the factor $iS$ in Eq. (2.3) can be transformed into S-S by an analytic continuation that leaves both the definition of the triangle area (2.1) and the integration measure (2.4) unchanged. This is achieved by rewriting the variables $\ell_{e}^{2}$ in terms of the $N_{e}$-dimensional spherical coordinates $(r,\Omega_{N_{e}-1})$ and then rotating the integration contour of the radial variable $r$ to the imaginary axis. For more details, please refer to App. B in our previous paper [29]. We may consider the following partition function instead of Eq. (2.3):
@@ 2.2 Spike and pinched geometry
In our previous paper [29], we numerically confirmed the suppression of pinched geometries through the finiteness of $\left<\tau^{2}\right>_{A}$ for a regular triangulation shown in Fig. 1. For different triangulations and integral measures, we verify the suppression by examining the higher moments $\left<\tau^{n}\right>_{A}$ up to $n=10$ in Sec. 3.
@@ 3.1 Discretization of tensor indices
Fig. 7 illustrates the transformation in Eq. (3.6) and the resulting network in Eq. (3.7) (the right panel). The edges of each triangle are discretized through the quadrature rule, and in Fig. 7 all of them are represented as contracted tensor indices. The transformation from the left to the right panel of Fig. 7 corresponds to summing over all internal indices associated with the space-like edges (blue links), resulting in a tensor network of rank-$4$ tensors. This operation can be carried out without any further truncation. Note that, while in Fig. 6 the integrals were associated with the shared links, after Fig. 7 these integrals are replaced by finite-dimensional summations. As will be seen later in Fig. 13, the same manipulation is applied there as well.
Due to this treatment, the value of the original network is, in principle, obtained by evaluating the network shown in Fig. 7 and then taking the limits $K\rightarrow\infty$ and $D\rightarrow\infty$. In the actual calculations presented later, we use sufficiently large values of $K_{s},K_{t}$ and $D$ such that the results are converged.
The expectation value of a local observable can also be evaluated using the tensor-network method. Let us consider four time-like edges (four red links) attached to a single space-like edge (blue link), and denote the corresponding variables by $x_{i},x_{j},x_{k},x_{l}$. We then consider an operator ${\cal O}(x_{i},x_{j},x_{k},x_{l})$ that depends only on these four links. The area $A$ and edge lengths $\ell_{e}$ are examples of operators belonging to this class. For operators of this type, the expectation value can be represented by a tensor network that contains a single impurity tensor. For the discretized network, such a tensor network is given as follows:
$I_{{\cal O}}={\rm Tr}\ \widetilde{T}_{x_{i}x_{j}x_{k}x_{l}}\prod T_{x_{m}x_{n}x_{o}x_{p}}\ ,\ \ \ \text{with}\ \ \ \widetilde{T}_{x_{i}x_{j}x_{k}k_{l}}={\cal O}(x_{i},x_{j},x_{k},x_{l})\ T_{x_{i}x_{j}x_{k}x_{l}}\ ,$ (3.8)
where $\widetilde{T}$ is the impurity tensor. The tensor networks $I_{{\cal O}}$ and $Z^{\prime}$ are independently evaluated using the tensor renormalization group. To obtain the expectation value $\langle{{\cal O}}\rangle$, one has to evaluate the ratio $I_{{\cal O}}/Z^{\prime}$ and then take the limits where $K_{s}$ and $K_{t}$ are sent to infinity.
@@ Acknowledgement
Data availability
The data that support the findings of this article are openly available [28].